# Mean-field limits for Consensus-Based Optimization and Sampling

### v3

@@ 1 Introduction
CBO and CBS methods are a novel family of algorithms providing an alternative to the popular particle swarm optimization (PSO) algorithm [Kennedy_Eberhart, Kennedy1997ThePS] that is found in many optimization toolboxes. Recent studies have shown that CBO is competitive with particle swarm optimization [totzeck2018numerical] for a number of benchmark functions, and able to solve high-dimensional optimization problems arising from machine learning [CBO-high-dim, MR4329816]. By now, these methods have also been adapted for various other tasks and application settings, including finance [MR4337731], rare event simulation [2023arXiv230409077A] and federated learning [CBO-federal-carrillo2023], to mention just a few. Compared to PSO however, this family of algorithms is amenable to theoretical analysis. Since current theoretical guarantees are only available on the level of the mean-field description, providing quantitative estimates quantifying the distance between the evolution of the interacting particles and the corresponding mean-field probability distribution is essential.
This introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
@@ 1.1 Consensus-Based Optimization (CBO)
Given an objective function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$, we consider the following version of Consensus-Based Optimization (CBO):
@@
$\mathcal{M}_{\beta}(\mu):=\frac{\int_{\mathbf{R}^{d}}x\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)},\qquad\mathcal{M}_{\beta}(\mu^{J}_{t}):=\frac{\sum_{j=1}^{J}X^{(j)}\operatorname{e}^{-\beta f\left(X^{(j)}_{t}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}_{t}\right)}}.$ (1.2)
@@
By design, CBO allows passage to the corresponding mean-field description. Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [PSO], it is possible to formally write down the corresponding mean-field equation:
@@
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ’current best’ that appears in PSO [CBO-PSO-zero-inertia-1]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical]. While the consensus point does not necessarily coincide with the global minimizer of the cost function $f$, [carrillo2018analytical, Theorem 4.2] shows that the distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large parameter $\beta$. The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
@@ 2 Main results
This section is organized as follows. In Subsection 2.1, we present the assumptions used throughout this work. Then, after stating well-posedness results for CBO and CBS in Subsection 2.2, we prove in Subsection 2.3 a key preparatory result for the mean-field limits. We then prove quantitative mean-field results for Consensus-Based Optimization and Sampling in Subsections 2.4 and 2.5, respectively.
@@ 2.1 Main assumptions
    There exist constants $c_{\ell},c_{u}>0$ and $u\geqslant\ell\geqslant 0$ such that
    $\forall x\in^{d},\qquad f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)},$ (2.2a)
    $\forall x\in^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.$ (2.2b)
@@
  • When $\ell=u=0$, assumption (H2) implies that $f$ is bounded from above.
  • Observe that (H1) implies (2.2a) with $u=s+1$.
@@
In the rest of this paper, we denote by $\mathcal{A}(s,\ell,u)$ the set of objective functions that satisfy Assumption H with these parameters. Our main results hold only for $f\in\mathcal{A}(s,\ell,\ell)$, i.e. under Assumption H with $\ell=u$. Many of the results we present hold only for values of $p$ in a range dictated by the interval of validity of the Wasserstein stability estimates presented in Subsection 3.1. In order to avoid repetitions, we introduce the associated notation
$p_{\mathcal{M}}(s,\ell):=\begin{cases}s+2,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$},\end{cases}\qquad p_{\mathcal{C}}(s,\ell):=\begin{cases}s+3,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$}.\end{cases}$ (2.3)
@@ 2.2 Well-posedness for CBO and CBS
In the case of CBO, Theorem 2.2 was already proved in [carrillo2018analytical, Theorem 2.1]. We now present well-posedness results for the mean-field SDEs, beginning with CBO. For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [carrillo2018analytical, Theorems 3.1 and 3.2].
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.3) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$. Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.3) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$, and it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.4)
@@
where $D(\overline{\rho},x):=S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}^{\mathsf{T}}$.
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$ and fix a final time $T>0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, there exists a strong solution $\overline{X}:\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.5) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.5) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, and it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}).$ (2.6)
@@ 2.3 Key preparatory lemma
We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of Markov’s inequality,
@@ 2.4 Quantitative mean-field limit for CBO
$b(x,\mu):=-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu):=S\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)},$ (2.10)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ has moments of sufficiently high order. Consider the systems (2.8) and (2.9) with the coefficients given by (2.10). Then, for each $p>0$, there exists a constant $C>0$ independent of $J$ such that
@@
We prove below the following more general statement: suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$. If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
@@
$W_{p}\Bigl{(}\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\Bigr{)}\leqslant CJ^{-\frac{1}{2}},$
@@
Here, $W_{p}$ denotes the pathwise Wasserstein-$p$ distance, associated with the following metric on $\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}$:
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,W_{p}(\mu,\nu).$
@@
    $\mathbf{E}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}:=\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
@@
where we used that the function $z\mapsto(q-z)/z^{2}$ is decreasing. Take $p\in[2\vee p_{\mathcal{M}},q/2]$ and fix $j\in\llbracket 1,J\rrbracket$. As in Sznitman’s argument [ReviewChaintronII, Theorem 3.1], we begin by writing
@@
$\Omega_{J,t}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\},$ (2.16)
where $R>\mathbf{E}[Z_{j}]$ is fixed. Define
$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}.$
By Theorem 2.3 it holds that $\mathbf{E}\left[\lvert Z_{j}\rvert^{q/p}\right]<\infty.$ Note that $q\geqslant 2p$. Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]$
@@
    $\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right)\leqslant 2^{q-1}\left(\kappa+\widetilde{\kappa}\right)$
    where $\widetilde{\kappa}:=C_{2}\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{q}\right]<\infty$. Here, $C_{2}>0$ is the constant from Proposition A.3. Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{\frac{(q-1)p}{q}}\left(\kappa+\widetilde{\kappa}\right)^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$

### v4

@@ 1 Introduction
CBO and CBS methods are a novel family of algorithms providing an alternative to the popular particle swarm optimization (PSO) algorithm [Kennedy_Eberhart, Kennedy1997ThePS] that is found in many optimization toolboxes. Recent studies have shown that CBO is competitive with PSO for a number of benchmark functions [totzeck2018numerical], and able to solve high-dimensional optimization problems arising from machine learning [CBO-high-dim, MR4329816]. For a detailed comparison between CBO and PSO, and additional context on the connections between the two methods, we refer the reader to the articles [CBO-PSO-zero-inertia-1, grassi2023mean, huang2023global].
By now, CBO and CBS have also been adapted for various other tasks and application settings, including finance [MR4337731], rare event simulation [2023arXiv230409077A] and federated learning [CBO-federal-carrillo2023], to mention just a few. Compared to PSO however, this family of algorithms is more amenable to theoretical analysis. Since current theoretical guarantees for CBO and CBS are only available at the level of the mean-field description, providing quantitative estimates for the distance between the evolution of the interacting particles and the corresponding mean-field probability distribution is essential.
The rest of this introductory section is organized as follows: we first introduce Consensus-Based Optimization (Subsection 1.1) and Consensus-Based Sampling (Subsection 1.2), giving pointers to the relevant literature. In Subsection 1.3 we present a brief review of the literature on the use of coupling methods to prove propagation of chaos for models with non-globally Lipschitz coefficients. Finally, we summarize the main contributions of this work in Subsection 1.4, and introduce the notation used throughout the paper in Subsection 1.5.
@@ 1.1 Consensus-Based Optimization (CBO)
Given an objective function $f\colon\mathbf{R}^{d}\rightarrow\mathbf{R}$ to be minimized, we consider the following version of Consensus-Based Optimization (CBO):
@@
$\mathcal{M}_{\beta}(\mu):=\frac{\int_{\mathbf{R}^{d}}x\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)},\qquad\mathcal{M}_{\beta}(\mu^{J}_{t}):=\frac{\sum_{j=1}^{J}X^{(j)}\operatorname{e}^{-\beta f\bigl{(}X^{(j)}_{t}\bigr{)}}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\bigl{(}X^{(k)}_{t}\bigr{)}}}.$ (1.2)
@@
By design, CBO allows passage to the corresponding mean-field description, which is not the case for the PSO algorithm. Indeed, since the particles in (1.1) evolve towards the weighted mean (1.2) instead of the current best particle, as is the case for Particle Swarm Optimization (PSO) [PSO], it is possible to formally write down the corresponding mean-field equation:
@@
where the weighted mean $\mathcal{M}_{\beta}(\mu)$ for general measures $\mu\in\mathcal{P}(\mathbf{R}^{d})$ can be interpreted as a smooth approximation of the ‘current best’ that appears in PSO [CBO-PSO-zero-inertia-1]. The connection between the particle system (1.1) and the dynamics (1.3), which is of McKean–Vlasov type and comparatively simpler to analyze mathematically, stimulated theoretical research using tools from stochastic analysis and partial differential equations [CBO-mfl-Huang2021, carrillo2018analytical, fornasier2021consensusbased]. In particular, consensus formation and convergence of the CBO mean-field dynamics (1.3) to a consensus point as $t\to\infty$ has been proven in [carrillo2018analytical]. While the consensus point does not necessarily coincide with the global minimizer of the cost function $f$, [carrillo2018analytical, Theorem 4.2] shows that the distance between the consensus point and the true global minimizer can be made arbitrarily small by choosing a large parameter $\beta$. The proof in [carrillo2018analytical] covers also non-convex cost functions $f$.
@@ 2 Main results
This section is organized as follows. In Subsection 2.1, we present the assumptions used throughout this work. Then, after stating well-posedness results for CBO and CBS in Subsection 2.2, we prove in Subsection 2.3 a key preparatory result for the mean-field limits. We then prove quantitative mean-field results for Consensus-Based Optimization and Sampling in Subsections 2.4 and 2.5, respectively. Finally, a numerical experiment illustrating the convergence as $J\to\infty$ for CBO is presented in Subsection 2.6.
@@ 2.1 Main assumptions
    There exist constants $c_{\ell},C_{\ell},c_{u},C_{u}>0$ and $u\geqslant\ell\geqslant 0$ such that
    $\forall x\in^{d},\qquad f(x)-f_{\star}\leqslant c_{u}|x|^{u}+C_{u},$ (2.2a)
    $\forall x\in^{d},\qquad f(x)-f_{\star}\geqslant c_{\ell}|x|^{\ell}-C_{\ell}.$ (2.2b)
In the rest of this paper, we denote by $\mathcal{A}(s,\ell,u)$ the set of objective functions that satisfy Assumption H with these parameters.
@@
  • The case $\ell=u=0$ corresponds to bounded $f$.
  • If $f\in\mathcal{A}(s,\ell,u)$, then necessarily $\ell\leqslant s+1$, because the set $\mathcal{A}(s,\ell,u)$ is empty for $\ell>s+1$.
@@
Our main results hold for $f\in\mathcal{A}(s,\ell,\ell)$, i.e. under Assumption H with $\ell=u$. The problem of proving mean-field limits for CBO/CBS in the case $\ell<u$ is discussed in Subsection 5.2.
Many of the results we present hold only for values of $p$ in a range dictated by the interval of validity of the Wasserstein stability estimates presented in Subsection 3.1. In order to avoid repetitions, we introduce the associated notation
$p_{\mathcal{M}}(s,\ell)\coloneq\begin{cases}s+2,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$},\end{cases}\qquad p_{\mathcal{C}}(s,\ell)\coloneq\begin{cases}s+3,\,&\text{if $\ell=0$},\\ 1,\,&\text{if $\ell>0$}.\end{cases}$ (2.3)
@@ 2.2 Well-posedness for CBO and CBS
In the case of CBO, Theorem 2.2 was already proved in [carrillo2018analytical, Theorem 2.1]. We now present well-posedness results for the mean-field SDEs, beginning with CBO. For $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$, the following result generalizes [carrillo2018analytical, Theorems 3.1 and 3.2].
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and fix a final time $T>0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, there exists a strong solution $\overline{X}\colon\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.3) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$. Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.3) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ is continuous over $[0,T]$, and it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}\coloneq\operatorname{Law}(\overline{X}_{t}).$ (2.4)
@@
where $D(\overline{\rho},x)\coloneq S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}S\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho})\bigr{)}^{\mathsf{T}}$.
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and fix a final time $T>0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, there exists a strong solution $\overline{X}\colon\Omega\to C([0,T],\mathbf{R}^{d})$ to (1.5) with initial condition $\overline{X}_{0}=x_{0}$ such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, Furthermore, the process $\overline{X}$ is unique within the class of strong solutions to (1.5) such that $t\mapsto\mathcal{M}_{\beta}(\overline{\rho}_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\overline{\rho}_{t})$ are continuous over $[0,T]$, and it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty,$ $\text{where}\qquad\overline{\rho}_{t}\coloneq\operatorname{Law}(\overline{X}_{t}).$ (2.6)
@@ 2.3 Key preparatory lemma
We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of Markov’s inequality.
@@ 2.4 Quantitative mean-field limit for CBO
$b(x,\mu)\coloneq-\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)}$ and $\sigma(x,\mu)\coloneq S\bigl{(}x-\mathcal{M}_{\beta}\left(\mu\right)\bigr{)},$ (2.10)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ has bounded moments of all orders. Consider the systems (2.8) and (2.9) with the coefficients given by (2.10). Then, for each $p>0$, there exists a constant $C>0$ independent of $J$ such that
@@
We prove below the following more general statement: Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ such that $\ell\leqslant s+1$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$. If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
@@
$\mathcal{W}_{p}\Bigl{(}\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\Bigr{)}\leqslant CJ^{-\frac{1}{2}},$
@@
Here, $\mathcal{W}_{p}$ denotes the pathwise Wasserstein-$p$ distance, associated with the following metric on $\Bigl{(}C\bigl{(}[0,T],\mathbf{R}^{d}\bigr{)}\Bigr{)}^{J}$:
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}\bigl{(}^{d}\bigr{)}\times\mathcal{P}_{p}\bigl{(}^{d}\bigr{)},\qquad\left\lvert\mathcal{M}_{\beta}(\mu)-\mathcal{M}_{\beta}(\nu)\right\rvert\leqslant\widetilde{L}_{\mathcal{M}}\,\mathcal{W}_{p}(\mu,\nu).$
@@
    $\mathbf{E}\left\lvert\mathcal{M}_{\beta}(\overline{\mu}^{J})-\mathcal{M}_{\beta}(\mu)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}},\qquad\overline{\mu}^{J}\coloneq\frac{1}{J}\sum_{j=1}^{J}\delta_{\overline{X}^{(j)}},\qquad\left\{\overline{X}^{(j)}\right\}_{j\in\mathbf{N}}\stackrel{{\scriptstyle\rm{i.i.d.}}}{{\sim}}\mu.$
@@
where we used that the function $z\mapsto(q-z)/z^{2}$ is decreasing. Take $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ and fix $j\in\llbracket 1,J\rrbracket$. As in Sznitman’s argument [ReviewChaintronII, Theorem 3.1], we begin by writing
@@
$\Omega_{J,t}\coloneq\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\},$ (2.16)
where
$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p},$
and $R>\mathbf{E}[Z_{j}]$ is fixed. By Theorem 2.3 it holds that $\mathbf{E}\left[\lvert Z_{j}\rvert^{q/p}\right]<\infty.$ Note that $q\geqslant 2p$. Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[\mathcal{W}_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]$
@@
    $\mathbf{E}\left[\mathcal{W}_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right)\leqslant 2^{q-1}\left(\kappa+\overline{\kappa}\right)$
    where $\overline{\kappa}:=C_{2}\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{q}\right]<\infty$. Here, $C_{2}>0$ is the constant from Proposition A.4. Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{\frac{(q-1)p}{q}}\left(\kappa+\overline{\kappa}\right)^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$